Accelerator Update for February 29, 2008 - March 3, 2008
Friday Feb 29
The day shift began with Operations monitoring the Tevatron (TeV) store 5943 (61.7E30) and the Recycler stash (269.3E10), with the Antiproton Source stacking (10.7E10), with the Main Injector (MI), Linac, and Booster operating normally, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams, with NuMI off for horn repair, and with the SciBooNE / MiniBooNE (BNB) experiments taking beam.
At 8:16 AM, Operations reported that a Linac quadrupole power supply’s (L:QPS111) output had sagged. Operators raised the output.
At 8:20 AM, two Booster vacuum pumps (B:IP61 & 72) tripped off. They reset.
A Linac expert began putting their multiwires into the beam at 11:02 AM to check beam positions.
MI experts began making MI-30 collimator measurements and adjustments at 11:14 AM.
At 11:30 AM, the Booster vacuum suddenly went bad. Experts investigated.
At 11:33 AM, Operations put Pbar into standby while the Booster vacuum problems were investigated.
The Neutron Therapy Facility (NTF) reported at 11:41 AM that they had finished treating their last patient for the day.
Operations reported that the Booster vacuum had recovered by 11:50 AM and that Booster had resumed sending beam to Pbar and BNB.
Operations terminated store 5943 at 1:12 PM.
Operations put the TeV into shot setup at 1:22 PM.
Booster experts reported at 1:39 PM that there was a vacuum leak in the BRF6 area.
At 2:02 PM, the 400MeV area tripped off for the first of two times in eight minutes due to bad beam positions. It reset.
Operations established store 5944 at 3:18 PM with an initial luminosity of 263.43E30.
Operations reported at 4 PM that the Recycler had stashed antiprotons four times during the day shift.
Operations reported that the TeV A and F sector front ends had been rebooted at 4:52 PM.
At 6:34 PM, the Pbar lithium lens tripped off. Operators contacted an expert and reset the lens.
At 6:41 PM, KRF5 tripped off. It reset.
At 6:48 PM, the lithium lens tripped off again. Operators contacted an expert who said that an access would be required. Operations began contacting people and preparing for a Pbar AP0 access.
At 7:53 PM, a Booster vacuum pump (IP61) tripped off. It reset.
At 8:19 PM, BRF10 tripped off. It reset.
Pbar experts began their access into AP0 at 9:31 PM.
At 10:22 PM, Pbar experts reported that the lens/transfer module needed to be replaced.
At 10:38 PM, KRF6 tripped off. It reset.
Operations reported at midnight that the Recycler had stashed two times during the evening shift.
Saturday March 1
The midnight shift began with Operations monitoring store 5944 (97.78E30) and the stash (112.3E10), with no stacking due to a lithium lens problem, with NuMI off for horn repair, and with BNB taking beam.
At 12:09 AM, BRF10 tripped off and would not reset. Operators bypassed the station and brought BRF19 on in its place.
At 1:11 AM, Operations reported that a problem with the TeV’s horizontal tune caused a loss in luminosity of about 3.0E30.
At 2:17 AM, the Booster East Anode Power Supply (EAPS) tripped off. Operators discovered that BRF5 had tripped off and would not reset. Operators bypassed the station and brought EAPS back online.
At 2:32 AM, a Booster vacuum pump (IP62) tripped off for the first of two times in two hours. It reset.
At 5:58 AM, Pbar experts reported that the 10 mm lens LCW circuit had failed its leak test. Since quite a few technicians had already been working for many hours on this repair, experts halted all work until later in the day when they would have a meeting to determine what to do next.
At 7:02 AM, BRF3 tripped off. It reset.
Operations began preparing Booster for an access at 7:19 AM (to repair the vacuum leak) by turning off beam to Booster.
At 8:37 AM, a Recycler kicker (KPS2A) alarmed on a magnet temperature indication. An operator contacted an expert.
Booster experts began their access at 9:07 AM.
At 9:11 AM, KRF6 tripped off. It reset.
A Booster vacuum expert reported at 10:53 AM that they had found a vacuum leak on an “O-ring” in the BRF6-1 area. They began repairs.
At 11:21 AM, KRF3 tripped off for the first of two times in about 80 minutes. It reset.
At 1:17 PM, KRF2 tripped off. It reset.
A Pbar expert accessed the AP0 vault area at 2:48 PM.
The Booster vacuum experts came out of access at 3:02 PM. The vacuum was recovering.
Operations established beam to BNB at 4:13 PM.
Operations halted beam to BNB temporarily, at 4:56 PM, to tune the Booster beam.
At 5:42 PM, a Booster vacuum pump (IP61) tripped off for the first of three times in two hours. It reset.
Operations reported at 6:10 PM that they were tuning on the MI-8 transfer line.
At 6:12 PM, LRF2 tripped off. It reset.
A MI expert began at study at 8:22 PM. This study halted beam to BNB.
At 10:38 PM, KRF3 tripped off. It reset.
The MI study ended at 11:08 PM and BNB resumed taking beam.
Sunday March 2
The midnight shift began with Operations monitoring store 5944 and the stash, with Pbar off for lens repair, with NuMI off for horn repair, and with BNB taking beam.
At 2:54 AM, KRF2 tripped off. It reset.
Operations terminated store 5944 at 3:02 AM.
At 3:36 AM, KRF 3 tripped off. It reset.
Operations put the TeV through a wet squeeze at 3:53 AM.
Operations put the TeV into shot setup at 5:36 AM.
At 6:40 AM, MIRF3 tripped off. It reset.
Operations began preparing MI for an access at 7:23 AM to inspect KPS2A.
Operations established store 5946 at 7:29 AM with an initial luminosity of 101.78E30.
At 7:34 AM, KRF3 tripped off. It reset.
At 8:29 AM, a Booster vacuum pump (IP61) tripped off. It reset.
A Booster expert began tuning the beam at 10:02 AM.
Recycler kicker experts reported at 10:26 AM that they found and fixed a small vapor leak and then filled the reservoir manifold with coolant.
Operations brought MI back on line at 12:06 PM and resumed sending beam to BNB.
Operations turned the Recycler on and injected reverse protons at 12:10 PM.
At 1:49 PM, Operations reported that the Linac momentum absorber vacuum was going into and out of alarm. An expert advised Operations to bypass the alarm for the time being. Experts would investigate the problem later.
A MI expert began investigating the beam coalescing at 3:51 PM.
At 6:49 PM, the Linac 400 MeV area tripped off for the first of four times in 24 minutes due to bad beam positions. It reset.
At 7:26 PM, KRF7 tripped off. It reset.
At 8:20 PM, the Linac 400 MeV area tripped off for the first of five times in 11 minutes. It reset. Operations tuned the H- Source.
The Pbar experts working at AP0 ran into further problems at 11:24 PM. They decided to stop working for the night.
Monday March 3
The midnight shift began with Operations monitoring store 5946 and the stash, with Pbar off due to lens work, with NuMI off for horn repair, and with BNB taking beam.
At 5:47 AM, KRF3 & 6 tripped off. They reset.
Operations terminated store 5946 at 7:10 AM.
The Plans for Wednesday
The plans for today are to repair the lithium lens and then resume stacking, stashing, and storing. During the lens repairs, TeV experts will use the time to also make repairs.
Comments and Suggestions
|last modified 03/03/2008 email Fermilab|
|Security, Privacy, Legal|